# LLMs Can’t Plan, But Can Help Planning in LLM-Modulo Frameworks

### 02 Feb 2024 v1

@@ 4 Related Work
We will note that while the LLM-Modulo framework is being proposed in general form here for the first time, there are certainly works in leveraging LLMs in planning and reasoning tasks that are in line with the spirit of the LLM-Modulo framework. For exmaple, both Valmeekam et al. (2023c) and Stechly et al. (2023) describe and evaluate a backprompting interaction between an LLM and an external verifier. Work on FunSearch Romera-Paredes et al. (2023) depends on a generate-test loop between a specially fine-tuned LLM that guesses solutions, and an external symbolic evaluator that critiques them. The authors note how the external verifier is critical for avoiding falling prey to hallucinations (i.e., approximate solution candidates that have flaws). AlphaGeometry Trinh et al. (2024) too depends on the Generate-Test-Critique interaction between a fine-tuned LLM and a symbolic evaluator. Both these systems fine-tune pre-trained LLMs with task specific synthetic data–the correctness of which is vetted with external simulators (as we discuss in 3.3).
@@
While we focused on PDDL planning tasks for the sake of concreteness, we believe that the essence of LLM-Modulo framework is equally applicable to other scenarios involving planning and reasoning–such as Reinforcement Learning with Simulators. Such RL systems rely on rewards as feedback to train a policy. Analogous to the plan evaluation and critiques performed by the respective critics in the LLM-Modulo framework, environment simulators provide the required feedback. This has particularly been useful but largely overlooked in cases where LLMs are used as an actor to generate an admissible plan by interacting with a simulator, for example in the case of AlfWorld Yao et al. (2023b); Shinn et al. (2023) and Minecraft Wang et al. (2023). As mentioned in Section 3, similar to extracting a domain model such as in the case of PDDL planning Guan et al. (2023), designing a reward model for the plan generation - feedback cycle is yet another potential use case that has been recently looked at for text-based Kwon et al. (2022); Hao et al. (2023) and robot manipulation Ma et al. (2023) domains.
@@
Interestingly, the fact that LLM’s can help come up with approximate quasi-symbolic transition models, reward models and models of high level actions has made a bigger splash in RL. This is because for far too long, researchers there have tried to spurn any high level models (lest that would involve depending on humans) and focused on learning to act from sensory information, under the name of “deep reinforcement learning.” Given the horrendous sample complexity of the DRL methods even in reaching a single subgoal, and the well known fact that even approximate symbolic models can help drastically improve the performance (c.f. Guan et al. (2022)), coupled with the fact that LLM’s are only too glad to dream up approximate models and goal recipes, there has been a performance revolution of sorts there Yao et al. (2023b); Liang et al. (2023); Wang et al. (2023). If we look beyond the improvements in these lower level goal seeking behaviors–especially in the presence of ergodic simulators, the RL approaches dependent on LLMs will encounter the same issues regarding subgoal interactions that our discussion of PDDL planning problems brought into focus. The LLM-Modulo inspired frameworks will thus, we believe, be equally relevant there. Indeed, SayCan Ahn et al. (2022) the earliest use of LLMs in generating policies in an RL-with-Simulator scenario, explicitly filters the action choices suggested by the LLM with the help of simulator.
@@ 5 Conclusion
This position paper is a modest attempt to combat both over-optimism and over-pessimism about the role of LLMs in planning and reasoning tasks. Our position is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in the so-called LLM-Modulo Frameworks in conjunction with external sound model-based verifiers. In support of this position, we summarized the literature questioning the claims about the planning and self-verification capabilities of LLMs by themselves. We also discussed how conflating approximate knowledge acquisition and generating executable plans of action are behind many of the claims about planning and verification abilities of LLMs. We then shared LLM-Modulo framework, our vision for a productive way to integrate the impressive idea generation/approximate knowledge provision capabilities of LLMs with external verifiers with correctness guarantees for robust planning. We discussed how planning in LLM-Modulo framework avoids inheriting the expressiveness and search-complexity issues of traditional symbolic planners, while retaining their soundness guarantees. As we discussed, LLM-Modulo frameworks are consistent with some of the most high-profile success stories of “neuro-symbolic” architectures, including AlphaGeometry and FunSearch.

### 06 Feb 2024 v2

@@ 4 Related Work
We will note that while the LLM-Modulo framework is being proposed in general form here for the first time, there are certainly works in leveraging LLMs in planning and reasoning tasks that are in line with the spirit of the LLM-Modulo framework. For exmaple, both Valmeekam et al. (2023c) and Stechly et al. (2023) describe and evaluate a backprompting interaction between an LLM and an external verifier. Work on FunSearch Romera-Paredes et al. (2023) depends on a generate-test loop between a specially fine-tuned LLM that guesses solutions, and an external symbolic evaluator that critiques them. The authors note how the external verifier is critical for avoiding falling prey to hallucinations (i.e., approximate solution candidates that have flaws). AlphaGeometry Trinh et al. (2024) too depends on the Generate-Test-Critique interaction between a fine-tuned LLM and a symbolic evaluator. Both these systems fine-tune pre-trained LLMs with task specific synthetic data–the correctness of which is vetted with external simulators (as we discuss in Section 3.3).
@@
While we focused on PDDL planning tasks for the sake of concreteness, we believe that the essence of LLM-Modulo framework is equally applicable to other scenarios involving planning and reasoning–such as Reinforcement Learning with Simulators. Such RL systems rely on rewards as feedback to train a policy. Simulators takes on the roles of plan evaluation and critiques performed by the respective critics in the LLM-Modulo framework (e.g. Rajvanshi et al. (2023)). The fact that simulators play the role of verifiers is often not explicitly recognized in cases where LLMs are used as an actor to generate an admissible plan by interacting with a simulator, for example in the case of AlfWorld Yao et al. (2023b); Shinn et al. (2023) and Minecraft Wang et al. (2023). As mentioned in Section 3, similar to extracting a domain model such as in the case of PDDL planning Guan et al. (2023), designing a reward model for the plan generation - feedback cycle is yet another potential use case that has been recently looked at for text-based Kwon et al. (2022); Hao et al. (2023) and robot manipulation Ma et al. (2023) domains.
@@
Interestingly, the fact that LLM’s can help come up with approximate quasi-symbolic transition models, reward models and models of high level actions has made a bigger splash in RL. This is because for far too long, researchers there have tried to spurn any high level models (lest that would involve depending on humans; Kambhampati (2021)) and focused on learning to act from sensory information, under the name of “deep reinforcement learning.” Given the horrendous sample complexity of the DRL methods even in reaching a single subgoal, and the well known fact that even approximate symbolic models can help drastically improve the performance (c.f. Guan et al. (2022)), coupled with the fact that LLM’s are only too glad to dream up approximate models and goal recipes, there has been a performance revolution of sorts there Yao et al. (2023b); Liang et al. (2023); Wang et al. (2023). If we look beyond the improvements in these lower level goal seeking behaviors–especially in the presence of ergodic simulators, the RL approaches dependent on LLMs will encounter the same issues regarding subgoal interactions that our discussion of PDDL planning problems brought into focus. The LLM-Modulo inspired frameworks will thus, we believe, be equally relevant there. Indeed, SayCan Ahn et al. (2022) the earliest use of LLMs in generating policies in an RL-with-Simulator scenario, explicitly filters the action choices suggested by the LLM with the help of simulator.
@@ 5 Conclusion
This position paper is a modest attempt to combat both over-optimism and over-pessimism about the role of LLMs in planning and reasoning tasks. Our position is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in the so-called LLM-Modulo Frameworks in conjunction with external sound model-based verifiers. In support of this position, we summarized the literature questioning the claims about the planning and self-verification capabilities of LLMs by themselves. We also discussed how conflating approximate knowledge acquisition and generating executable plans of action is behind many of the claims about planning and verification abilities of LLMs. We then shared LLM-Modulo framework, our vision for a productive way to integrate the impressive idea generation/approximate knowledge provision capabilities of LLMs with external verifiers with correctness guarantees for robust and expressive planning. We discussed how planning in LLM-Modulo framework avoids inheriting the expressiveness and search-complexity limitations of traditional symbolic planners, while retaining their soundness guarantees. As we discussed, LLM-Modulo frameworks are consistent with some of the most high-profile success stories of “neuro-symbolic” architectures, including AlphaGeometry and FunSearch.
Acknowledgments
The ideas discussed in this paper have evolved over a series of talks, tutorials and twitter threads. The discussions, feedback and encouragement from colleagues, including Daniel Borrajo, Tom Dietterich, Yann LeCun, Sarath Sreedharan, and Dan Weld is gratefully acknowledged.